# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Field Theory Perspective

### v4

@@ 1.1 Activation Functions
An activation function is a function that acts on the output of each layer of a neural network. The remarkable success of deep neural networks is closely related to the choice of appropriate nonlinear activation functions, and there has been extensive discussion solely on the research of activation functions. Early studies on neural networks include the perceptron (rosenblatt58), where the activation function is the sign function, and it was composed of only one or two layers. The perceptron had the drawback that its derivative was almost everywhere zero, making it difficult to apply the backpropagation algorithm. According to (cybenko89), for a neural network to satisfy the universal approximation property, the activation function must be bounded, non-constant, monotonically increasing, and continuous. In the 1980s and 1990s, smooth activations like the sigmoid activation and tanh activation, along with their variations, were proposed (Rumelhart86, Narayan97, LeCun02). These activations had the disadvantage of gradient vanishing when stacking layers deeply. To overcome this drawback, the Rectified Linear Unit (ReLU) was proposed by (Nair10), whose derivative is a non-zero constant. Variations of ReLU, such as PReLU, GELU, and ELU, have also been proposed in the 2010s (He15, Hendrycks16, Clervert15). Recently, there have been attempts to find activations tailored to specific tasks to impart inductive bias to the neural network structure according to the task (Ramachandran17, Sun24).
@@ 1.2 Rectified Power Unit
One of the attempts to provide good regularity and inductive bias to the architecture is the Rectified Power Unit (RePU), a generalization of ReLU (Li20). Early results using RePU include the application of ReQU, a type of RePU, in the DeepRitz method, which is a neural network approach for solving Partial Differential Equations (PDEs) (E18). RePU has also been proposed to construct differentiable neural networks, and error bounds related to this have been analyzed (Li20). According to research, using a RePU network allows for the optimal approximation of smooth functions in Sobolev space with optimal depth and width, and the optimal approximation error bound is also provided in Li20. Similarly, another study derived that a shallow RePU network can approximate multivariate functions of a certain function class with finite neural network weight norms and possibly unbounded width (Abdeljawad22). Furthermore, it was shown that the derivatives of a network composed of RePU can be expressed as a mixture of various types of RePU, and an improved upper bound on the complexity of these derivatives was derived. This demonstrates the capability of RePU by showing the risk bound for the deep score matching estimator (DSME) (Shen23). Despite these results highlighting the good regularity of RePU neural networks, experimental evidence shows that RePU networks suffer from exploding or vanishing values and fail to train properly when the layers are deep.
@@ 1.3 Effective Field Theory of Neural Networks
The language of quantum field theory (QFT) is useful for analyzing systems with a very high degree of freedom. One such example is artificial neural networks. It is known that the distribution of neural network ensembles follows a Gaussian process when the width is infinite, corresponding to the free field in the QFT framework (Neal96, Rasmussen04, Roberts22, Banta24, Halverson21). In practice, neural networks with finite width exhibit non-Gaussianity in their distributions as layers are stacked, allowing the distribution of neurons to be calculated perturbatively, as in weakly-interacting theory (Banta24, Halverson21). Interestingly, when the width of the neural network is finite, the propagation of information through the neural network can be understood as a renormalization group (RG) flow. By analyzing the recursive relations of the kernels in RG flow, hyperparameters can be tuned to a criticality state that ensures the stability of the network (Roberts22). Additionally, this analysis allows for the classification of activation functions into universality classes based on the behavior of the RG flow (Roberts22). Some types of activation functions, according to this classification, cannot satisfy criticality conditions from the outset.
@@ 2.1 Overview of Neural Networks and Activation Functions
Sigmoid $\sigma(z)=\frac{1}{1+e^{-z}}$ cybenko89
Tanh $\sigma(z)=\tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ cybenko89, LeCun02
Sin $\sigma(z)=\sin(z)$ Parascandolo16, Sitzmann20
@@
Softplus $\sigma(z)=\log(1+e^{z})$ Bishop06
SWISH $\sigma(z)=\frac{z}{1+e^{-z}}$ Ramachandran17
GELU $\sigma(z)=\Big[\frac{1}{2}+\frac{1}{2}\text{erf}\Big(\frac{z}{\sqrt{2}}\Big)\Big]z$ Hendrycks16
@@
Definition 3 (Rectified Power Unit (RePU). The RePU activation function is defined by the following equation:
@@ 2.2 Effective Field Theory for Neural Networks
$\begin{split}\mathbb{E}[z_{i;\alpha}^{(1)}]&=\mathbb{E}\bigg[b_{i}^{(1)}+\sum_{j=0}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}\bigg]=0,\\ \mathbb{E}[z_{i_{1};\alpha_{1}}^{(1)}z_{i_{2};\alpha_{2}}^{(1)}]&=\mathbb{E}\bigg[\bigg(b_{i_{1}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{1}j_{1}}^{(1)}x_{j_{1};\alpha_{1}}\bigg)\bigg(b_{i_{2}}^{(1)}+\sum_{j=0}^{n_{0}}W_{i_{2}j_{2}}^{(1)}x_{j_{2};\alpha_{2}}\bigg)\bigg]\\ &=\delta_{i_{1}i_{2}}\bigg(C_{b}^{(1)}+C_{W}^{(1)}\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}x_{j;\alpha_{1}}x_{j;\alpha_{2}}\bigg)=\delta_{i_{1}i_{2}}G_{\alpha_{1}\alpha_{2}}^{(1)}.\end{split}$
@@
It is known that the following relationship holds for the coefficients. This can be derived by comparing the coefficients of the connected correlator given by the mean metric and four-point vertex with those of the connected correlator assumed by the ansatz provided by action $S$. $g^{\alpha_{1}\alpha_{2}}_{(l+1)}$ and $v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l+1)}$ (Roberts22):
@@
Kernels and Representation Group flow Now, we will investigate how the statistics of the preactivation change as they pass through the layers. We will refer to the evolution of the distribution through the layers as the representation group flow, following the terminology of (Roberts22). While it is not feasible to explicitly calculate the statistics for a complex non-Gaussian distribution, it is possible to develop a perturbative theory to understand the recursive flow of the key statistics of the distribution. Before delving into this, let’s define the notation for Gaussian expectations. For an observable $\mathcal{O}(z)$ with a mean-zero Gaussian distribution with variance $\bf{K}=(K_{ij})$, the expectation is denoted as $\langle\mathcal{O}(z)\rangle_{\bf{K}}$. Now, if we define the kernel of the neural network through the following recursive relation:
@@
it is known that this kernel satisfies following properties, as discussed in (Roberts22).
$\begin{split}G^{(l)}_{\alpha_{1}\alpha_{2}}&=K^{(l)}_{\alpha_{1}\alpha_{2}}+O\Big(\frac{1}{n_{l-1}}\Big),\\ V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}&=\Big(C_{W}^{(l+1)}\Big)^{2}\Big[\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &-\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}-\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &+\frac{n_{l}}{4n_{l-1}}\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}\Big\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\Big(z_{\beta_{1}}z_{\beta_{2}}-K_{\beta_{1}\beta_{2}}^{(l)}\Big)\Big\rangle_{K^{(l)}}\\ &\times\Big\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\Big(z_{\beta_{3}}z_{\beta_{4}}-K_{\beta_{3}\beta_{4}}^{(l)}\Big)\Big\rangle_{K^{(l)}}\Big]+O\Big(\frac{1}{n_{l}}\Big).\end{split}$ (5)
@@
Then, it is known that the following recursion can be derived (Roberts22):
@@ 3.1 Susceptibility Calculation
Activation functions that exhibit similar behaviors according to the RG flow are classified into universality classes. Universality class is a concept in statistical physics used to classify systems that exhibit similar behavior at the limit scale. According to (Roberts22), scale-invariant activations (e.g. ReLU, LeakyReLU) form a single universality class called the self-invariant universality class. Tanh and Sin fall into the $K^{\star}=0$ universality class, while SWISH and GELU belong to the half-stable universality class. On the other hand, Perceptron, Sigmoid, and Softplus are known to not satisfy criticality at all. We will now calculate the susceptibility for the RePU activation function following Eq. (6) and demonstrate that RePU also fails to satisfy criticality. For the RePU activation function of order $p$, the parallel susceptibility and perpendicular susceptibility are calculated as follows:
@@ 5 Approximation Properties of MRePU networks
In this section, we theoretically demonstrate that a neural network constructed with the MRePU activation function can reconstruct polynomials approximately, similar to RePU neural networks, and achieves the same universal approximation property for the class of differentiable functions as shown in (Shen23). Additionally, we experimentally verify that, beyond this theoretical foundation, neural networks with the MRePU activation function possess an appropriate inductive bias when approximating certain classes of functions.
@@ 5.1 Approximation Theorems for MRePU
In (Shen23), it is shown that a RePU network of $p$-th order can accurately represent an expanded univariate polynomial using Horner’s method, utilizing $4p$ nodes for multiplication and $2p$ nodes to construct a $p$-th order polynomial. Furthermore, by applying these univariate polynomial results inductively, the representation of multivariate polynomials can also be achieved. Therefore, for the MRePU of order $p$ network, if we can demonstrate that it can approximate any $p$-th order polynomial using $2p$ nodes, as in the case of RePU, we obtain the same result. The following lemma formalizes this.
@@
Lemma 10. A polynomial of degree $p$, $\sum_{k=0}^{p}a_{k}x^{k}$ can be represented by a $p$-th order MRePU network with one hidden layer and at most $2p$ hidden nodes.
Proof Consider the neural network constructed as follows.
$y=W_{1}^{T}\sigma_{m;p}(W_{0}^{T}x+b_{0})+v_{0}$
Here, each parameter is defined as follows:
$W_{0}=\begin{bmatrix}1\\ -1\\ \vdots\\ 1\\ -1\end{bmatrix}\in\mathbb{R}^{2p},\quad b_{0}=\begin{bmatrix}u_{1}-1\\ -u_{1}-1\\ \vdots\\ u_{p}-1\\ -u_{p}-1\end{bmatrix}\in\mathbb{R}^{2p},\quad W_{1}=\begin{bmatrix}v_{1}\\ (-1)^{p}v_{1}\\ \vdots\\ v_{p}\\ (-1)^{p}v_{p}\end{bmatrix}\in\mathbb{R}^{2p}.$
where $u_{1},\dots,u_{p}$ are pairwise distinct real numbers and $v_{0},v_{1},\dots,v_{p}$ are trainable weights. For each pair $(x+u_{k}-1)$ and $(-x-u_{k}-1)$ passed through the $p$-th order MRePU,
$\begin{split}&v_{k}(x+u_{k})^{p}(x+u_{k}-1)+v_{k}(x+u_{k})^{p}(-x-u_{k}-1)\\ &=-2v_{k}(x+u_{k})^{p}.\end{split}$
Summing over $k$ gives
$\begin{split}y(x)&=-2\sum_{k=1}^{p}v_{k}(x+u_{k})^{p}+v_{0}\\ &=-2\sum_{k=1}^{p}\sum_{j=0}^{p}\binom{p}{j}u_{k}^{p-j}x^{j}+v_{0}.\end{split}$
Comparing coefficients with $\sum_{j=0}^{p}a_{j}x^{j}$ yields, for $j=1,\dots,p$,
$2\sum_{l=1}^{p}\binom{p}{k}u_{l}^{p-k}v_{l}=a_{k}.$
This can be written as the linear system:
$\begin{bmatrix}1&\dots&1\\ \vdots&&\vdots\\ u_{1}^{p-2}&\dots&u_{p}^{p-2}\\ u_{1}^{p-1}&\dots&u_{p}^{p-1}\par\end{bmatrix}\begin{bmatrix}v_{1}\\ \vdots\\ v_{p-1}\\ v_{p}\end{bmatrix}=-\frac{1}{2}\begin{bmatrix}a_{p}\binom{p}{p}^{-1}\\ \vdots\\ a_{2}\binom{p}{2}^{-1}\\ a_{1}\binom{p}{1}^{-1}\end{bmatrix}$. (12)
Since $u_{1},\dots,u_{p}$ are distinct, $V(u)$ is invertible, and hence $v_{1},\dots,v_{p}$ are uniquely determined. Finally, matching the constant term $v_{0}=a_{0}+2\sum_{k=1}^{p}v_{k}u_{k}^{p}$, any polynomial of degree $p$ can be represented by a single-hidden-layer $p$-th order MRePU network with at most $2p$ hidden nodes. $\square$
@@
Then, by replacing Lemma 40, which is used in the proof of Theorem 3 in (Shen23), we obtain the following theorem:
@@
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=12pN^{d-1}+6p(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1}),\\ &\mathcal{U}=(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}),\\ &\mathcal{P}=(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}).\end{split}$
Additionally, as a corollary to the above theorem, we obtain a universal approximation result for the class of differentiable functions, which corresponds to Theorem 5 in (Shen23). Before stating this result, we define a reference norm that will serve as our standard.
@@
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=12pN^{d-1}+6p(N^{d-1}-N)/(N-1),\\ &\mathcal{U}=(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1),\\ &\mathcal{P}=(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1).\end{split}$
such that for $|\alpha|\leq\min\{s,N\}$, we have
$sup_{K}|D^{\alpha}(f-\psi_{N})|\leq C_{p,s,d,K}\|f\|_{C^{|\alpha|}}N^{-(s-|\alpha|)},$

### v5

@@ 1.1 Activation Functions
An activation function is a function that acts on the output of each layer of a neural network. The remarkable success of deep neural networks is closely related to the choice of appropriate nonlinear activation functions, and there has been extensive discussion solely on the research of activation functions. Early studies on neural networks include the perceptron (Rosenblatt (1958)), where the activation function is the sign function, and it was composed of only one or two layers. The perceptron had the drawback that its derivative was almost everywhere zero, making it difficult to apply the backpropagation algorithm. According to (Cybenko (1989)), for a neural network to satisfy the universal approximation property, the activation function must be bounded, non-constant, monotonically increasing, and continuous. In the 1980s and 1990s, smooth activations like the sigmoid activation and tanh activation, along with their variations, were proposed (Rumelhart et al. (1986), Narayan (1997), LeCun et al. (2002)). These activations had the disadvantage of gradient vanishing when stacking layers deeply. To overcome this drawback, the Rectified Linear Unit (ReLU) was proposed by (Nair and Hinton (2010)), whose derivative is a non-zero constant. Variations of ReLU, such as PReLU, GELU, and ELU, have also been proposed in the 2010s (He et al. (2015), Hendrycks and Gimpel (2016), Clevert et al. (2015)). Recently, there have been attempts to find activations tailored to specific tasks to impart inductive bias to the neural network structure according to the task (Ramachandran et al. (2017), Sun et al. (2024)).
@@ 1.2 Rectified Power Unit
One of the attempts to provide good regularity and inductive bias to the architecture is the Rectified Power Unit (RePU), a generalization of ReLU (Li et al. (2020)). Early results using RePU include the application of ReQU, a type of RePU, in the DeepRitz method, which is a neural network approach for solving Partial Differential Equations (PDEs) (E and Yu (2018)). RePU has also been proposed to construct differentiable neural networks, and error bounds related to this have been analyzed (Li et al. (2020)). According to research, using a RePU network allows for the optimal approximation of smooth functions in Sobolev space with optimal depth and width, and the optimal approximation error bound is also provided in Li et al. (2020). Similarly, another study derived that a shallow RePU network can approximate multivariate functions of a certain function class with finite neural network weight norms and possibly unbounded width (Abdeljawad and Grohs (2022)). Furthermore, it was shown that the derivatives of a network composed of RePU can be expressed as a mixture of various types of RePU, and an improved upper bound on the complexity of these derivatives was derived. This demonstrates the capability of RePU by showing the risk bound for the deep score matching estimator (DSME) (Shen et al. (2023)). Despite these results highlighting the good regularity of RePU neural networks, experimental evidence shows that RePU networks suffer from exploding or vanishing values and fail to train properly when the layers are deep.
@@ 1.3 Effective Field Theory of Neural Networks
The language of quantum field theory (QFT) is useful for analyzing systems with a very high degree of freedom. One such example is artificial neural networks. It is known that the distribution of neural network ensembles follows a Gaussian process when the width is infinite, corresponding to the free field in the QFT framework (Neal (1996), Rasmussen (2004), Roberts et al. (2022), Banta et al. (2024), Halverson et al. (2021)). In practice, neural networks with finite width exhibit non-Gaussianity in their distributions as layers are stacked, allowing the distribution of neurons to be calculated perturbatively, as in weakly-interacting theory (Banta et al. (2024), Halverson et al. (2021)). Interestingly, when the width of the neural network is finite, the propagation of information through the neural network can be understood as a renormalization group (RG) flow. By analyzing the recursive relations of the kernels in RG flow, hyperparameters can be tuned to a criticality state that ensures the stability of the network (Roberts et al. (2022)). Additionally, this analysis allows for the classification of activation functions into universality classes based on the behavior of the RG flow (Roberts et al. (2022)). Some types of activation functions, according to this classification, cannot satisfy criticality conditions from the outset.
@@ 2.1 Overview of Neural Networks and Activation Functions
Sigmoid $\sigma(z)=\frac{1}{1+e^{-z}}$ Cybenko (1989)
Tanh $\sigma(z)=\tanh(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$ Cybenko (1989), LeCun et al. (2002)
Sin $\sigma(z)=\sin(z)$ Parascandolo et al. (2016), Sitzmann et al. (2020)
@@
Softplus $\sigma(z)=\log(1+e^{z})$ Bishop (2006)
SWISH $\sigma(z)=\frac{z}{1+e^{-z}}$ Ramachandran et al. (2017)
GELU $\sigma(z)=\Big[\frac{1}{2}+\frac{1}{2}\text{erf}\Big(\frac{z}{\sqrt{2}}\Big)\Big]z$ Hendrycks and Gimpel (2016)
@@
Definition 3 (Rectified Power Unit (RePU)). The RePU activation function is defined by the following equation:
@@ 2.2 Effective Field Theory for Neural Networks
$\begin{split}\mathbb{E}[z_{i;\alpha}^{(1)}]&=\mathbb{E}\bigg[b_{i}^{(1)}+\sum_{j=1}^{n_{0}}W_{ij}^{(1)}x_{j;\alpha}\bigg]=0,\\ \mathbb{E}[z_{i_{1};\alpha_{1}}^{(1)}z_{i_{2};\alpha_{2}}^{(1)}]&=\mathbb{E}\bigg[\bigg(b_{i_{1}}^{(1)}+\sum_{j=1}^{n_{0}}W_{i_{1}j_{1}}^{(1)}x_{j_{1};\alpha_{1}}\bigg)\bigg(b_{i_{2}}^{(1)}+\sum_{j=1}^{n_{0}}W_{i_{2}j_{2}}^{(1)}x_{j_{2};\alpha_{2}}\bigg)\bigg]\\ &=\delta_{i_{1}i_{2}}\bigg(C_{b}^{(1)}+C_{W}^{(1)}\frac{1}{n_{0}}\sum_{j=1}^{n_{0}}x_{j;\alpha_{1}}x_{j;\alpha_{2}}\bigg)=\delta_{i_{1}i_{2}}G_{\alpha_{1}\alpha_{2}}^{(1)}.\end{split}$
@@
It is known that the following relationship holds for the coefficients. This can be derived by comparing the coefficients of the connected correlator given by the mean metric and four-point vertex with those of the connected correlator assumed by the ansatz provided by action $S$. $g^{\alpha_{1}\alpha_{2}}_{(l+1)}$ and $v^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l+1)}$ (Roberts et al. (2022)):
@@
Kernels and Representation Group flow Now, we will investigate how the statistics of the preactivation change as they pass through the layers. We will refer to the evolution of the distribution through the layers as the representation group flow, following the terminology of (Roberts et al. (2022)). While it is not feasible to explicitly calculate the statistics for a complex non-Gaussian distribution, it is possible to develop a perturbative theory to understand the recursive flow of the key statistics of the distribution. Before delving into this, let’s define the notation for Gaussian expectations. For an observable $\mathcal{O}(z)$ with a mean-zero Gaussian distribution with variance $\bf{K}=(K_{ij})$, the expectation is denoted as $\langle\mathcal{O}(z)\rangle_{\bf{K}}$. Now, if we define the kernel of the neural network through the following recursive relation:
@@
it is known that this kernel satisfies following properties, as discussed in (Roberts et al. (2022)).
$\begin{split}G^{(l)}_{\alpha_{1}\alpha_{2}}&=K^{(l)}_{\alpha_{1}\alpha_{2}}+O\Big(\frac{1}{n_{l-1}}\Big),\\ V^{(l+1)}_{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}&=\Big(C_{W}^{(l+1)}\Big)^{2}\Big[\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\\ &-\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\rangle_{K^{(l)}}\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\rangle_{K^{(l)}}\Big]\\ &+\frac{n_{l}}{4n_{l-1}}\sum_{\beta_{1},...,\beta_{4}\in\mathcal{D}}V^{(\alpha_{1}\alpha_{2})(\alpha_{3}\alpha_{4})}_{(l)}\Big\langle\sigma(z_{\alpha_{1}})\sigma(z_{\alpha_{2}})\Big(z_{\beta_{1}}z_{\beta_{2}}-K_{\beta_{1}\beta_{2}}^{(l)}\Big)\Big\rangle_{K^{(l)}}\\ &\times\Big\langle\sigma(z_{\alpha_{3}})\sigma(z_{\alpha_{4}})\Big(z_{\beta_{3}}z_{\beta_{4}}-K_{\beta_{3}\beta_{4}}^{(l)}\Big)\Big\rangle_{K^{(l)}}+O\Big(\frac{1}{n_{l}}\Big).\end{split}$ (5)
@@
Then, it is known that the following recursion can be derived (Roberts et al. (2022)):
@@ 3.1 Susceptibility Calculation
Activation functions that exhibit similar behaviors according to the RG flow are classified into universality classes. Universality class is a concept in statistical physics used to classify systems that exhibit similar behavior at the limit scale. According to (Roberts et al. (2022)), scale-invariant activations (e.g. ReLU, LeakyReLU) form a single universality class called the self-invariant universality class. Tanh and Sin fall into the $K^{\star}=0$ universality class, while SWISH and GELU belong to the half-stable universality class. On the other hand, Perceptron, Sigmoid, and Softplus are known to not satisfy criticality at all. We will now calculate the susceptibility for the RePU activation function following Eq. (6) and demonstrate that RePU also fails to satisfy criticality. For the RePU activation function of order $p$, the parallel susceptibility and perpendicular susceptibility are calculated as follows:
@@ 5 Approximation Properties of MRePU networks
In this section, we provide a theoretical justification that neural networks equipped with the MRePU activation can reconstruct polynomials, in close analogy to RePU networks, and consequently inherit the same universal approximation property for the class of differentiable functions established in Shen et al. (2023). Beyond this theoretical guarantee, we empirically demonstrate that MRePU networks exhibit a suitable inductive bias when approximating certain function classes. Furthermore, we show through experiments on standard benchmarks such as MNIST and CIFAR-10 that, under the proposed criticality condition, training remains stable and reliably converges.
@@ 5.1 Approximation Theorems for MRePU
In (Shen et al. (2023)), it is shown that a RePU network of $p$-th order can accurately represent an expanded univariate polynomial using Horner’s method, utilizing $4p$ nodes for multiplication and $2p$ nodes to construct a $p$-th order polynomial. Furthermore, by applying these univariate polynomial results inductively, the representation of multivariate polynomials can also be achieved. Therefore, for the MRePU of order $p$ network, if we can demonstrate that it can construct any $p$-th order polynomial using $\mathcal{O}(p)$ nodes, as in the case of RePU, we obtain the same result. The following lemma formalizes this.
@@
Lemma 10. A polynomial of degree $p$, $P(x)=\sum_{k=0}^{p}a_{k}x^{k}$ can be represented by a $p$-th order MRePU network with one hidden layer and at most $4(p+1)$ hidden nodes.
Proof The proof proceeds in three steps.
Step 1: A 4-unit exact block for $(x-c)^{p}$ on $\mathbb{R}$. Fix any $c\in\mathbb{R}$ and write t:=xct:=x-c. Consider the four units
$\sigma_{m;p}(t-1),\qquad\sigma_{m;p}(2t-1),\qquad\sigma_{m;p}(-t-1),\qquad\sigma_{m;p}(-2t-1).$
We first compute them on the two half-lines.
Case 1: $t\geq 0$. Then $t-1\geq-1$ and $2t-1\geq-1$, while t1<1-t-1<-1 and 2t1<1-2t-1<-1. Hence
$\sigma_{m;p}(t-1)=(t-1)t^{p}=t^{p+1}-t^{p},\qquad\sigma_{m;p}(2t-1)=(2t-1)(2t)^{p}=2^{p+1}t^{p+1}-2^{p}t^{p},$
and $\sigma_{m;p}(-t-1)=\sigma_{m;p}(-2t-1)=0$. Choose
$A:=-2,\qquad B:=2^{-p}.$
Then
$A\sigma_{m;p}(t-1)+B\sigma_{m;p}(2t-1)$ $=A(t^{p+1}-t^{p})+B(2^{p+1}t^{p+1}-2^{p}t^{p})$
$=(A+B2^{p+1})t^{p+1}+(-A-B2^{p})t^{p}$
$=\bigl(-2+2^{-p}2^{p+1}\bigr)t^{p+1}+\bigl(2-2^{-p}2^{p}\bigr)t^{p}$
$=0\cdot t^{p+1}+1\cdot t^{p}$
$=t^{p}.$
Case 2: $t\leq 0$. Then t1<1t-1<-1 and 2t1<12t-1<-1, while $t-1\geq-1$ and $2t-1\geq-1$. Thus $\sigma_{m;p}(t-1)=\sigma_{m;p}(2t-1)=0$ and
$\sigma_{m;p}(-t-1)=(-t-1)(-t)^{p}=(-t)^{p+1}-(-t)^{p},$
$\sigma_{m;p}(-2t-1)=(-2t-1)(-2t)^{p}=2^{p+1}(-t)^{p+1}-2^{p}(-t)^{p}.$
With the same $A,B$ as above,
$A\sigma_{m;p}(-t-1)+B\sigma_{m;p}(-2t-1)$ $=A\bigl((-t)^{p+1}-(-t)^{p}\bigr)+B\bigl(2^{p+1}(-t)^{p+1}-2^{p}(-t)^{p}\bigr)$
$=(A+B2^{p+1})(-t)^{p+1}+(-A-B2^{p})(-t)^{p}$
$=0\cdot(-t)^{p+1}+1\cdot(-t)^{p}$
$=(-t)^{p}.$
Since $(-t)^{p}=(-1)^{p}t^{p}$, multiplying by $(-1)^{p}$ yields
$(-1)^{p}\bigl(A\sigma_{m;p}(-t-1)+B\sigma_{m;p}(-2t-1)\bigr)=t^{p}.$
Combining the two cases, for every $t\in\mathbb{R}$ we have the identity
$t^{p}=A\sigma_{m;p}(t-1)+B\sigma_{m;p}(2t-1)+(-1)^{p}\Bigl(A\sigma_{m;p}(-t-1)+B\sigma_{m;p}(-2t-1)\Bigr),$ (12)
with A=2A=-2 and $B=2^{-p}$. Substituting t=xct=x-c gives an exact 4-unit one-hidden-layer $\sigma_{m;p}$-network representation of $(x-c)^{p}$ on $\mathbb{R}$.
Step 2: Spanning order $p$ polynomials by shifted $p$-th powers. Choose distinct real numbers $c_{0},c_{1},\dots,c_{p}$ (e.g. $c_{i}=i$). For each $i$,
$(x-c_{i})^{p}=\sum_{j=0}^{p}\binom{p}{j}x^{j}(-c_{i})^{p-j}.$
We claim that for any coefficients $(a_{0},\dots,a_{p})$ there exist $(\alpha_{0},\dots,\alpha_{p})$ such that
$P(x)=\sum_{i=0}^{p}\alpha_{i}(x-c_{i})^{p}.$ (13)
Indeed, comparing coefficients of $x^{j}$ in (13) gives the linear system
$\sum_{i=0}^{p}\alpha_{i}(-c_{i})^{p-j}=\frac{a_{j}}{\binom{p}{j}},\qquad j=0,1,\dots,p.$
The matrix $V=[(-c_{i})^{p-j}]_{j,i}$ is a Vandermonde matrix up to a reversal of rows, and since the $c_{i}$ are distinct it is invertible. Hence the system admits a (unique) solution $(\alpha_{0},\dots,\alpha_{p})$, proving (13).
Step 3: Assemble the network for $P$. For each $i\in\{0,\dots,p\}$, represent $(x-c_{i})^{p}$ using the 4-unit identity (12) (with $t=x-c_{i}$), and multiply the four corresponding output weights by $\alpha_{i}$. Summing these $(p+1)$ blocks yields a single one-hidden-layer $\sigma_{m;p}$ network with $4(p+1)$ hidden units whose output equals $\sum_{i=0}^{p}\alpha_{i}(x-c_{i})^{p}=P(x)$ for all $x\in\mathbb{R}$.
Therefore any polynomial of degree at most $p$ is represented exactly on $\mathbb{R}$ by a one-hidden-layer $p$-MRePU network with at most $4(p+1)$ units. $\square$
@@
Then, by replacing Lemma 40, which is used in the proof of Theorem 3 in (Shen et al. (2023)), we obtain the following theorem:
@@
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=24(p+1)N^{d-1}+12(p+1)(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1}),\\ &\mathcal{U}=(12p+14)(2N^{d}-N^{d-1}-N)+(4p+4)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}),\\ &\mathcal{P}=(60p+62)(2N^{d}-N^{d-1}-N)+(4p+5)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}).\end{split}$
Additionally, as a corollary to the above theorem, we obtain a universal approximation result for the class of differentiable functions, which corresponds to Theorem 5 in (Shen et al. (2023)). Before stating this result, we define a reference norm that will serve as our standard.
@@
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=24(p+1)N^{d-1}+12(p+1)(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1}),\\ &\mathcal{U}=(12p+14)(2N^{d}-N^{d-1}-N)+(4p+4)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}),\\ &\mathcal{P}=(60p+62)(2N^{d}-N^{d-1}-N)+(4p+5)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}).\end{split}$